a body is thrown vertically upwards and takes 5 seconds to reach maximum height . the distance travelled by the body will be same in

Grade:11

2 Answers

raosuresh623
4 years ago
V=u+gt, here at top v=0,  g is downward and a is upward ie. a=-g. 0= u-gt 
u=gt 
u=10×5
u=50m/s
For distance 
V^2-u^2=-2gh
0-2500=-2×10×s
s=125m
That is answer
 
Arun
4 years ago
Dear student
 
1st and 10th second is correct answer.
using g = 10 m/sec^2 : 
Vo = g*t = 10*5 = 50 m/sec 

from t = 0 to t = 1 sec 
h = ((Vo*1-5*1^2)-0) = (50-5) = 45 m (up) 

from t = 9 to t = 10 sec 
h = ((Vo*10-5*10^2)-(Vo*9-5*9^2) 
h= (0-450+405) = -45 m (down)